OpenStudy (anonymous):

Multiply (u-5)(u+5) simplify your answer u^2-5u-5u-25?

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

very close check the sign of one of the middle ones :)

OpenStudy (anonymous):

is it u^2+5u-5u-25?

OpenStudy (sleepyhead314):

yes! then just add like terms :)

OpenStudy (anonymous):

u^2-25? do the 2 5u cancel each other?

OpenStudy (sleepyhead314):

yep! u^2 - 25 is correct ^_^ this is a rule you can keep in mind for similar problems if you want (a + b)(a - b) = a^2 - b^2

OpenStudy (anonymous):

awesome thank you :-) I have more also im sorry

OpenStudy (sleepyhead314):

it's alright :) I'm glad I can help you :D

OpenStudy (academicgurusinc):

Alright so this problem is talking about the difference of squares. Think of (a+b)(a-b) = a^2-b^2 .... here we have the similar case. So (u-5)(u+5)=u^2-25 as our final answer. If you found this helpful, I encourage to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)